OpenStudy (anonymous):

Bethany has $3.10 in nickels and dimes. She has a total of 44 coins. What is the value of the nickels Bethany has? a. $0.90 b. $1.30 c. $1.80 d. $2.60

OpenStudy (anonymous):

Okay I will help you @justsmile531 k just give me a minute to solve

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I say A

OpenStudy (anonymous):

I am not sure though

OpenStudy (texaschic101):

n + d = 44 ----> d = 44 - n 0.05n + .10d = 3.10 now sub 44 - n in for d in the 2nd equation 0.05n + .10(44 - n) = 3.10 0.05n + 4.40 - 0.10n = 3.10 0.05n - 0.10n = 3.10 - 4.40 -0.05n = - 1.30 n = -1.30/-0.05 n = 26 n + d = 44 26 + d = 44 d = 44 - 26 d = 18 check.. 0.05n + 0.10d = 3.10 0.05(26) + 0.10(18) = 3.10 1.30 + 1.80 = 3.10 3.10 = 3.10 (correct) There is 26 nickels and 18 dimes. And if she has 26 nickels, she has $ 1.30 in nickels

OpenStudy (texaschic101):

@justsmile531 ...your answer is b

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

sure thing :)